# Neutral phylogenetic models and their role in tree-based biodiversity measures

### v2

@@ 1 Introduction
In this paper we deal with both fixed phylogenetic trees and random birth-death trees. A phylogenetic tree $T$ here refers to a rooted directed tree (having its edges directed away from the root) with a set of labelled leaves (vertices of out-degree 00) and with the remaining interior vertices unlabelled and of out-degree at least 2. If $X$ is the set of leaves of $T$ we say that $T$ is a phylogenetic $X$–tree. A star tree is a phylogenetic tree with a single interior vertex. Mostly we deal with binary trees in which all interior vertices have out-degree exactly 2. Although we can view a phylogenetic tree as a (discrete) graph without edge lengths, we will often also assign a (positive real-valued) length $\ell_{e}$ to each edge $e$ of the tree (and if these edge lengths are proportional to time they are said to be ultrametric). If we remove the labels on the leaves of a binary phylogenetic tree (and remove any edge lengths) we obtain a tree shape denoted $\tau$. For example, when $n=4$, there are precisely two tree shapes: the balanced (‘fork’) shape and the unbalanced (‘caterpillar’) shape.
A birth–death tree $\mathcal{T}_{t}$ is a rooted tree generated by a stochastic process of binary splitting and death of lineages, starting from a single lineage at time 00 and acting for time $t$. The tree $\mathcal{T}_{t}$ is complete if it includes lineages that become extinct by time $t$, whereas a reduced birth-death tree is obtained from the complete tree by removing these extinct lineages (this reduced tree is also referred to in the some literature as the reconstructed tree). Notice that a reduced birth-death tree with extant species set $X$ induces a discrete phylogenetic $X$–tree (by ignoring the edge lengths, and the ‘stem’ edge at the top of the tree), as indicated in Fig. 1(lower left). For either class of trees, a pendant edge in a tree is an edge incident with a leaf; all other edges are interior.
@@
By convention, a vertex of out-degree 2 in a phylogenetic tree represents an ancestral speciation event. This would typically involves a lineage splitting off from an existing lineage and forming a new species. However, there is usually no distinction made in a birth-death (or phylogenetic) tree as to which of the two edges corresponds to the formation of a ‘new’ species, rather than a continuation of the existing one. This distinction has been noted by others (e.g. [4, 5]) and it can change how one measures quantities, such as the number of times new species were generated in the path that leads to a given present-day species. To address this, we use the term evolutionary tree for a tree that makes explicit which of the two edges directed out of a degree-2 vertex corresponds to the newly forming species (the horizontal edges in Fig. 1 upper (right)). Evolutionary trees also show lineages that lead to extinct taxa, and so we define the reduced evolutionary tree to be the tree obtained from an evolutionary tree by (i) ignoring extinct lineages, and (ii) not treating a lineage as a new species if it branches off a lineage that has no surviving taxa at the present. This is illustrated in Fig. 1 (the two trees on the lower right). Once again, ignoring edge lengths in a reduced evolutionary tree leads to a discrete phylogenetic tree. These concepts are illustrated in Fig. 1 (the circled vertices in the figure correspond to the root of the discrete phylogenetic tree).
Notation
Given a rooted phylogenetic $X$–tree $T$, an edge $e$ of $T$, and a leaf $x$ of $T$ we let:
  • $C_{T}(e)$ be the set of leaves of $T$ that are descendants of edge $e$;
  • $n_{e}=|C_{T}(e)|$;
  • $P_{T}(x)$ be the set of edges on the path from the root of $T$ to leaf $x$.
@@ 2.1 Depth of a random leaf in reduced trees
Notice that for the reduced birth-death tree (or the phylogenetic tree) in Fig. 1 the depth $D_{3}$ of a randomly chosen leaf equals 2 with probability $2/3$ and equals $1$ with probability $1/3$. By contrast, in the reduced evolutionary trees in this figure there are two (equally probable) scenarios to consider. In the first (the left-hand reduced evolutionary tree in Fig. 1, the depth of a randomly chosen leaf is uniform on 0,1,2; in the other the depth is 0 with probability $1/3$ and 1 with probability $2/3$ (the right-hand reduced evolutionary tree in Fig. 1). Thus $D^{\prime}_{3}=0,1,2$ with probabilities: $\frac{1}{3},\frac{1}{2}$ and $\frac{1}{6}$, respectively.
@@
Firstly, if $T_{n}$ is a YH tree, and we select uniformly at random one of the two subtrees incident with the root of $T_{n}$ then the number $N$ of leaves in this subtree tree is uniformly distributed between $1$ and $n-1$ (this follows by a simple Pólya-Urn model argument). By contrast, the PDA model leads to a quite different distribution for $N$, since the probability that one of the two subtrees incident with the root consists of just one leaf equals $n/(2n-3)$ which converges to $1/2$, as $n\rightarrow\infty$ (whereas the corresponding probability converges to 00 under the YH model).
@@
Next, suppose that a pair of distinct leaves of $T_{n}$ (say, $x,y$) is selected uniformly at random, and consider the number $N_{xy}$ of edges that lie between the root of $T_{n}$ and the most recent common ancestor of $x$ and $y$. In that case, the distribution of $N_{xy}$ converges to a geometric distribution with a mean of $2$ [16]. In particular, the probability that the most recent common ancestor of $x$ and $y$ coincides with the root of $T_{n}$ converges to $\frac{1}{3}$. This last claim is a special result of an earlier result by Sanderson [17]: for a randomly selected subset of leaves $S$ of size $k$, the probability that its most recent common ancestor of $S$ is precisely the root of $T_{n}$ converges to $1-\frac{2}{k+1}$ as $n\rightarrow\infty$.
@@
This claim follows by a slight extension of the argument used to establish Proposition 5 in [18]. Moreover, since every rooted tree shape is a subtree of $\tau_{k}$ for some (sufficiently large) $k$ it follows that the probability that a Yule tree $T_{n}$ contains a subtree of shape $\tau$ tends to $1$ as $n\rightarrow\infty$. Now, this last result also holds for the PDA model, but the proof is more involved (for details see [18]) since Claim 1 (above) fails to hold for the PDA model.
@@ 2.3 Edge lengths of Yule trees
Next, consider the Yule process, which has birth rate $\lambda$ and death rate $\mu=0$. In this model, each lineage splits according to an exponential distribution with rate $\lambda$. Suppose we sample a Yule tree just before the number of leaves changes from $n$ to $n+1$.
Suppose we select an edge uniformly at random from either (i) the set of pendant edges or (ii) the set of interior edges. In either case, the length of the sampled edge is exponentially distributed with expected value $\frac{1}{2\lambda}$. Thus the notion that the pendant edges might tend to be shorter, on average, than the interior edges (because the pendant edges are in some sense ‘still growing’) is misleading. For further details, and extensions to birth-death trees see [19]. By contrast, if we select one of the two edges descended from the root of this Yule tree, then the expected length of this edge is different; it is precisely $\frac{1}{\lambda}\left(1-\frac{1}{n}\right)$ [20].
We can also consider the sum of the lengths of all the edges of a Yule tree (the ‘phylogenetic diversity’ of all the species at the tips, considered further in Section 3). Let $S_{n}$ denote this sum at the last moment when the tree has $n\geq 2$ leaves. Since there are $22n-$ 2 edges in a binary tree with $n$ leaves, and since a randomly selected edge has expected length $1/2\lambda$, it follows that $S_{n}$ has expected value $(n-1)/\lambda$. However, more can be said: $S_{n}$ has a Gamma distribution with variance $(n-1)/\lambda^{2}$, and is asymptotically normally distributed as $n\rightarrow\infty$ [20].
@@ 2.4 The longest pendant and shortest interior edges in birth-death trees
Conditional on $N_{t}>0$, the ratio $L_{t}/t$ converges in probability to $\frac{1}{2}$ and
$\mathbb{E}\left[\frac{L_{t}}{t}|N_{t}>0\right]\to\frac{1}{2}$
as $\lambda t\rightarrow\infty$.
@@ 3.1 PD indices
PD provides a measure of diversity for subsets of species. However, conservation biologists are also interested in quantifying how much each species contributes to the total PD of a phylogenetic $X$-tree $T$. In other words, we are interested in functions $I_{(T,\ell)}:X\rightarrow\mathbb{R}^{\geq 0}$ for which
@@ 3.2 The loss of PD under extinction at the present
Mass extinction events have played an important role in evolution [34] and we are currently entering a new (anthropogenic) one. The simplest process of extinction at the present is the field-of-bullets (FOB) model, in which species become extinct independently, and with the same probability $\epsilon$. A more nuanced process – the generalised field-of-bullets model (gFOB) – assigns each species its own extinction probability $\epsilon_{x}$ while still assuming independence. In practice, species extinction probabilities can be estimated from their Red List status (from the International Union for Conservation of Nature) [35, 36]. More realistic extinction processes would account for the fact that the extinction of some species may trigger the further extinction of other species (‘extinction cascades’); alternatively, species that share common features may be correlated in their extinction risk if these features turn out to be deleterious. We do not consider these more complex models here.
@@ 3.3 Loss of PD on birth–death trees under extinction at the present
Let $\mathcal{T}_{t}$ be a birth–death tree with $\lambda\geq\mu$, and apply the simple FOB model, with survival probability $s$, to the extant leaves of $\mathcal{T}_{t}$. What proportion of the tree remains? More precisely, let $X_{t}$ denote the extant leaves of $\mathcal{T}_{t}$ and let $Y_{t}$ denote the subset of $X_{t}$ of the species that survive the FOB extinction event. What then can be said about the proportion of PD that survives, or, in other words, the ratio $PD_{\mathcal{T}_{t}}(Y_{t})/PD_{\mathcal{T}_{t}}(X_{t})$? As $t$ grows, it can be shown that this ratio converges in probability to a constant $\varphi_{PD}(s)$; moreover, $\varphi_{PD}(s)=\lim_{t\rightarrow\infty}\frac{\mathbb{E}[Y_{t}]}{\mathbb{E}[X_{t}]}$. When $\mu=0$ (i.e., a Yule tree), $\varphi_{PD}$ has a particularly simple form, namely:
@@
In all cases, $\varphi_{PD}(s)$ is a monotonically increasing concave function that satisfies $\varphi_{PD}(s)\geq s$ with equality if and only if $s\in\{0,1\}$. This is shown in Fig. 3, which also illustrates an interesting feature of $\varphi_{PD}(s)$: as $\mu$ approaches $\lambda$ from below, the function $y=\varphi_{PD}(s)$ approaches the unit step function (which is zero at $s=0$ and is $1$ for all other values of $s$). This is because the resulting critical trees, conditional on non-extinction, have very short pendant and near-pendant edges.
@@
Moreover, for any value of $\mu$ and $\lambda$, as the survival probability ($s$) approaches 1 the function $y=\varphi_{PD}(s)$ has a slope equal to the average pendant edge length divided by the average edge length. The intuition behind this is that when $s$ is close to 1, the only edges that disappear are isolated pendant edges (interior edges only start to disappear if two adjacent leaves disappear). Thus, for Yule trees, the slope of $y=\varphi_{PD}(s)$ at $s=1$ is $\frac{1}{2}$, whereas for birth–death trees with $\mu>0$, the slope at $s=1$ is smaller (for further details, see [38]).
Heuristic derivation of $\varphi_{PD}$: To provide some intuition for eqn. (6), here we give a brief (but non-rigorous) derivation of it. First recall that constant-rate birth-death trees induce a tree shape that follows the YH distribution, and for this distribution, the expected number of edges that have $k\leq n-1$ leaves as descendants is $2n/(k(k+1))$. Next, observe that the length of each such edge contributes to the PD score with probability $(1-(1-s)^{k})$ (i.e. iff one of these $k$ leaves survives). Now for a Yule tree, the expected length of the (interior and pendant) edges is the same, and so $\varphi_{PD}(s)$ reduces to $\sim\sum_{k=1}^{n}\frac{1}{k(k+1)}(1-(1-s)^{k})$ which converges to $s\ln(1/s)/(1-s)$ as $n\rightarrow\infty$.
@@ 3.4 Stochastic diversity indices for extinction at the present
Let $X$ denote a set of extant species, and consider the (random) subset $\mathcal{S}$ of species in $X$ that will still be extant (i.e. still present) at time $t_{*}>t$ under some extinction model. Let $(T,\ell)$ be an associated phylogenetic tree with leaf set $X$. Focusing on any particular species $x\in X$, let $\mathcal{S}_{x}=\mathcal{S}\setminus\{x\}$ and let
@@
Thus $\psi_{x}$ measures the expected increase in PD at time $t_{*}$ if species $x$ is present compared with if it is absent at time $t_{*}$ (noting that other species may also be extinct at time $t_{*})$. If we now specialise the extinction model to the gFOB model, where each species $s$ in $X$ has probability $\epsilon_{s}$ of being extinct at time $t_{*}$, then:
@@
where $\mathcal{I}_{x}\in\{\emptyset,\{x\}\}$, and $\epsilon_{x}=\mathbb{P}(\mathcal{I}_{x}=\emptyset)$. Thus, $\psi^{\prime}_{x}$ is the expected increase in PD at time $t_{*}$ if species $x$ is present at time $t_{*}$ compared with when its state is uncertain at time $t_{*}$. Straightforward algebra leads to the following relationship between $\psi_{x}$ and $\psi^{\prime}_{x}$ under the gFOB model:

### v3

@@ 1 Introduction
In this paper we deal with both fixed phylogenetic trees and random birth-death trees. A phylogenetic tree $T$ here refers to a rooted directed tree (having its edges directed away from the root) with a set of labelled leaves (vertices of out-degree 00) and with the remaining interior vertices unlabelled and of out-degree at least 2. If $X$ is the set of leaves of $T$ we say that $T$ is a phylogenetic $X$–tree. A star tree is a phylogenetic tree with a single interior vertex. Mostly we deal with binary trees in which all interior vertices have out-degree exactly 2. Although we can view a phylogenetic tree as a (discrete) graph without edge lengths, we will often also assign a (positive real-valued) length $\ell_{e}$ to each edge $e$ of the tree (and if these edge lengths are proportional to time they are said to be ultrametric). If we remove the labels on the leaves of a binary phylogenetic tree (and remove any edge lengths) we obtain a tree shape denoted $\tau$. For example, when $|X|=4$, there are precisely two tree shapes: the balanced (‘fork’) shape and the unbalanced (‘caterpillar’) shape.
Notation: Given a rooted phylogenetic $X$–tree $T$, an edge $e$ of $T$, and a leaf $x$ of $T$ we let: $n=|X|$; the number of leaves of $T$; $C_{T}(e)$ denote the set of leaves of $T$ that are descendants of edge $e$; $n_{e}=|C_{T}(e)|$; and $P_{T}(x)$ denote the set of edges on the path from the root of $T$ to leaf $x$.
A birth–death tree $\mathcal{T}_{t}$ is a rooted tree generated by a stochastic process of binary splitting and death of lineages, starting from a single lineage at time 00 and acting for time $t$. The tree $\mathcal{T}_{t}$ is complete if it includes lineages that become extinct by time $t$, whereas a reduced birth-death tree is obtained from the complete tree by removing these extinct lineages (this reduced tree is also referred to in the literature as the reconstructed tree). Notice that a reduced birth-death tree with extant species set $X$ induces a discrete phylogenetic $X$–tree (by ignoring the edge lengths, and the ‘stem’ edge at the top of the tree), as indicated in Fig. 1(lower left). For either class of trees, a pendant edge in a tree is an edge incident with a leaf; all other edges are interior.
@@
By convention, a vertex of out-degree 2 in a phylogenetic tree represents an ancestral speciation event. This would typically involves a lineage splitting off from an existing lineage and forming a new species. However, there is usually no distinction made in a birth-death (or phylogenetic) tree as to which of the two edges corresponds to the formation of a ‘new’ species, rather than a continuation of the existing one. This distinction has been noted by others (e.g. [4, 5], and invoking the asymmetry between parent and offspring lineages (as Yule did in his original model of genera evolution) can change how one measures various quantities. To address this, we use the term evolutionary tree for a tree that makes explicit which of the two edges directed out of a degree-2 vertex corresponds to the newly forming species (the horizontal edges in Fig. 1 upper (right)).
Evolutionary trees also show lineages that lead to extinct taxa, and so we define the reduced evolutionary tree to be the tree obtained from an evolutionary tree by (i) ignoring extinct lineages, and (ii) not treating a lineage as a new species if it branches off a lineage that has no surviving taxa at the present. This is illustrated in Fig. 1 (the two trees on the lower right). Once again, ignoring edge lengths in a reduced evolutionary tree leads to a discrete phylogenetic tree, as illustrated in Fig. 1 (the circled vertices in the figure correspond to the root of the discrete phylogenetic tree).
@@ 2.1 Depth of a random leaf in reduced trees
Notice that for the reduced birth-death tree (or the phylogenetic tree) in Fig. 1 the depth $D_{3}$ of a randomly chosen leaf equals 2 with probability $2/3$ and equals $1$ with probability $1/3$. By contrast, in the reduced evolutionary trees in this figure there are two scenarios to consider. In the first (the left-hand reduced evolutionary tree in Fig. 1), the depth of a randomly chosen leaf is uniform on 0,1,2; in the other (the right-hand reduced evolutionary tree in Fig. 1) the depth is 0 with probability $1/3$ and 1 with probability $2/3$ (the right-hand reduced evolutionary tree in Fig. 1). Since the two scenarios are equally likely it follows that $D^{\prime}_{3}=0,1,2$ with probabilities $\frac{1}{3},\frac{1}{2}$ and $\frac{1}{6}$, respectively.
@@
Firstly, if $T_{n}$ is a YH tree, and we select uniformly at random one of the two subtrees incident with the degree-2 root of $T_{n}$ then the number $N$ of leaves in this subtree tree is uniformly distributed between $1$ and $n-1$ (this follows by a simple Pólya-Urn model argument). By contrast, the PDA model leads to a quite different distribution for $N$, since the probability that one of the two subtrees incident with the root consists of just one leaf equals $n/(2n-3)$ which converges to $1/2$, as $n\rightarrow\infty$ (whereas the corresponding probability converges to 00 under the YH model).
@@
Next, given a YH tree with $n$ leaves, if we select uniformly at random a pair of leaves, then their most recent common ancestor (MRCA) vertex is likely to be ‘close’ to the root vertex of the tree. More precisely, the number of edges between the MRCA and the root converges in distribution (as $n$ grows) to a geometric random variable $G$ with $\mathbb{P}(G=g)=\frac{1}{3}\cdot(\frac{2}{3})^{g}$ for $g=0,1,\ldots$ [16]. Thus, the probability that the MRCA vertex of the chosen pair of leaves is the root of the YH tree converges to $\frac{1}{3}$ as $n$ grows  [^2] More generally, Sanderson [17] showed that for a randomly selected subset of leaves $S$ of size $k$, the probability that the MRCA of $S$ is the root of a YH tree converges to $1-\frac{2}{k+1}$ as $n\rightarrow\infty$.. For the PDA model however, the probability that the MRCA of the chosen pair of leaves is the root of the tree tends to zero as $n\rightarrow\infty$.
@@
This claim follows by a slight extension of the argument used to establish Proposition 5 in [18]. Moreover, since every rooted tree shape is a subtree of $\tau_{k}$ for some (sufficiently large) $k$ it follows that the probability that a YH tree $T_{n}$ contains a subtree of shape $\tau$ tends to $1$ as $n\rightarrow\infty$. This last result also holds for the PDA model, but the proof is more involved (for details see [18]) since Claim 1 (above) is easily shown to fail for the PDA model.
@@ 2.3 Edge lengths of Yule trees
Next, consider the Yule process, a continuous-time Markov process on state space $1,2,3\ldots,$ with a linear birth rate $\lambda$ and a death rate of 00. Such a process corresponds to the growing number of lineages in a random rooted binary tree in which each lineage of the tree splits independently according to an exponential distribution with rate $\lambda$. This tree is commonly referred to as Yule tree and we adopt this terminology here, despite it being somewhat contrived (as noted by Amaury Lambert in this special issue).
Suppose we now sample a Yule tree just before the number of leaves of the tree changes from $n$ to $n+1$. Ignoring the initial stem edge, select any other edge of this tree uniformly at random from either (i) the set of pendant edges or (ii) the set of interior edges. In either case, the length of the sampled edge is exponentially distributed with expected value $\frac{1}{2\lambda}$. Thus the notion that the pendant edges might tend to be shorter, on average, than the interior edges (because the pendant edges are in some sense ‘still growing’) is misleading. For further details, and extensions to birth-death trees see [19]. By contrast, if we select one of the two edges descended from the root of this Yule tree, then the expected length of this edge is different; it is precisely $\frac{1}{\lambda}\left(1-\frac{1}{n}\right)$ [20].
We can also consider the sum of the lengths of all the edges of a Yule tree (again ignoring the stem edge). Let $S_{n}$ denote this sum at the last moment when the tree has $n\geq 2$ leaves. Since there are $22n-$ 2 edges in a binary tree with $n$ leaves, and since a randomly selected edge has expected length $1/2\lambda$, it follows that $S_{n}$ has expected value $(n-1)/\lambda$. However, more can be said: $S_{n}$ has a Gamma distribution with variance $(n-1)/\lambda^{2}$, and is asymptotically normally distributed as $n\rightarrow\infty$ [20].
@@ 2.4 The longest pendant and shortest interior edges in birth-death trees
Conditional on $N_{t}>0$, the ratio $L_{t}/t$ converges in probability to $\frac{1}{2}$ and $\mathbb{E}\left[\frac{L_{t}}{t}|N_{t}>0\right]\to\frac{1}{2}$ as $\lambda t\rightarrow\infty$.
@@ 3.1 PD indices
Phylogenetic diversity provides a measure of diversity for subsets of species. However, conservation biologists are also interested in quantifying how much each species contributes to the total PD of a phylogenetic $X$-tree $T$. In other words, they are interested in functions $I_{(T,\ell)}:X\rightarrow\mathbb{R}^{\geq 0}$ for which
@@ 3.2 The loss of PD under extinction at the present
Mass extinction events have played an important role in evolution [34] and we are currently entering a new (anthropogenic) one. The simplest process of extinction at the present is the field-of-bullets (FOB) model, in which species become extinct independently, and with the same probability $\epsilon$. A more nuanced process – the generalised field-of-bullets model (gFOB) – assigns each species its own extinction probability $\epsilon_{x}$ while still assuming independence. In practice, species extinction probabilities can be estimated from their Red List status (from the International Union for Conservation of Nature) [35, 36]. More realistic extinction processes would account for the fact that the extinction of some species may trigger the further extinction of other species (leading to ‘extinction cascades’); alternatively, species that share common features may be correlated in their extinction risk if these features turn out to be deleterious. We do not consider these more complex models here.
@@ 3.3 Loss of PD on birth–death trees under extinction at the present
Let $\mathcal{T}_{t}$ be a birth–death tree with $\lambda\geq\mu$, and apply the simple FOB model, with survival probability $s$, to the extant leaves of $\mathcal{T}_{t}$. What proportion of the tree remains? More precisely, let $X_{t}$ denote the extant leaves of $\mathcal{T}_{t}$ and let $Y_{t}$ denote the subset of $X_{t}$ of the species that survive the FOB extinction event. What then can be said about the proportion of PD that survives, or, in other words, the ratio $PD_{\mathcal{T}_{t}}(Y_{t})/PD_{\mathcal{T}_{t}}(X_{t})$? From [38] the following convergence in probability holds as $t$ becomes large:
[…]
where
$\varphi_{PD}(s):=\lim_{t\rightarrow\infty}\frac{\mathbb{E}[PD_{\mathcal{T}_{t}}(Y_{t})]}{\mathbb{E}[PD_{\mathcal{T}_{t}}(X_{t})]}.$
When $\mu=0$ (i.e., a Yule tree), $\varphi_{PD}(s)$ has a particularly simple form, namely:
@@
In all cases, $\varphi_{PD}(s)$ is a monotonically increasing concave function that satisfies $\varphi_{PD}(s)\geq s$ with equality if and only if $s\in\{0,1\}$. This is shown in Fig. 3, which also illustrates an interesting feature of $\varphi_{PD}(s)$: as $\mu$ approaches $\lambda$ from below, the function $y=\varphi_{PD}(s)$ approaches the unit step function (which is zero at $s=0$ and is $1$ for $s\in(0,1]$). Roughly speaking, this is because the resulting critical trees, conditional on non-extinction, have very short pendant and near-pendant edges.
@@
Moreover, for any value of $\mu$ and $\lambda$, as the survival probability ($s$) approaches 1 the function $y=\varphi_{PD}(s)$ has a slope equal to the expected proportion of the PD of the tree that is contributed by its pendant edges. The intuition behind this is that when $s$ is close to 1, the only edges that disappear are isolated pendant edges (interior edges only start to disappear when two adjacent leaves disappear). Thus, for Yule trees, the slope of $y=\varphi_{PD}(s)$ at $s=1$ is $\frac{1}{2}$, whereas for birth–death trees with $\mu>0$, the slope at $s=1$ is smaller.
Heuristic derivation of $\varphi_{PD}$: To provide some intuition for eqn. (6), here we give a brief (but non-rigorous) derivation of it. First recall that constant-rate birth-death trees induce a tree shape that follows the YH distribution, and for this distribution, the expected number of edges that have $k\leq n-1$ leaves as descendants is $2n/(k(k+1))$. Next, observe that the length of each such edge contributes to the PD score with probability $(1-(1-s)^{k})$ (i.e. iff at least one of these $k$ leaves survives). Now for a Yule tree, the expected length of the (interior and pendant) edges is the same, and so $\varphi_{PD}(s)$ reduces to $\sim\sum_{k=1}^{n}\frac{1}{k(k+1)}(1-(1-s)^{k})$ which converges to $s\ln(1/s)/(1-s)$ as $n\rightarrow\infty$.
@@ 3.4 Stochastic diversity indices for extinction at the present
Let $X$ denote a set of extant species, and consider the (random) subset $\mathcal{S}$ of species in $X$ that will still be extant (i.e. still present) at some future time $t_{*}>t$ under some extinction model. Let $(T,\ell)$ be an associated phylogenetic tree with leaf set $X$. Focusing on any particular species $x\in X$, let $\mathcal{S}_{x}=\mathcal{S}\setminus\{x\}$ and let
@@
Thus $\psi_{x}$ measures the expected additional PD at time $t_{*}$ if species $x$ is present compared with if it is absent at time $t_{*}$ (noting that other species may also be extinct at time $t_{*})$. If we now specialise the extinction model to the gFOB model, where each species $s$ in $X$ has probability $\epsilon_{s}$ of being extinct at time $t_{*}$, then:
@@
where $\mathcal{I}_{x}\in\{\emptyset,\{x\}\}$, and $\epsilon_{x}=\mathbb{P}(\mathcal{I}_{x}=\emptyset)$. Thus, $\psi^{\prime}_{x}$ is the expected additional PD at time $t_{*}$ if species $x$ is present at time $t_{*}$ compared with if its state is uncertain at time $t_{*}$. Straightforward algebra leads to the following relationship between $\psi_{x}$ and $\psi^{\prime}_{x}$ under the gFOB model: